An airliner flies against the wind on one leg of its round trip in 5.5 hours. It flies back to its starting point, with the wind in 5 hours. it the average speed of the wind is 16 miles per hour, what is the speed of the airliner in still air?

Answer by mananth(16949)

An airliner flies against the wind on one leg of its round trip in 5.5 hours.
let speed of plane in still air be x
speed against wind = x-16
time = 5.5 hours
It flies back to its starting point, with the wind in 5 hours.
speed with wind = x+16
time = 5 hours
distance is same
d = rt


(x-16)*5.5 = (x+16)*5
(x-16)*11/2 = (x+16)*5
11x-176 = 10x+160
x= 176+160
x=336 mph